Sally can paint a room in 8 hours while it takes steve 9 hours to the same room. How long would it take them to paint the room if they worked together?

type an integer or a decimal round to one decimal place as needed


1
Expert's answer
2022-04-04T14:18:32-0400

"Sally: \frac{1}{8}\,\,room\,\,in\,\,an\,\,hour\\Steve: \frac{1}{9}\,\,room\,\,in\,\,an\,\,hour\\Together: \frac{1}{8}+\frac{1}{9}=\frac{17}{72}\,\,room\,\,in\,\,an\,\,hour\\\frac{1}{\frac{17}{72}}=\frac{72}{17}\,\,hours\,\,for\,\,a\,\,room"